OpenStudy (mony01):

Find the area of the largest rectangle that can be inscribed in a right triangle with legs of lengths 3 cm and 8 cm if two sides of the rectangle lie along the legs.

OpenStudy (dumbcow):

 find x,y which maximize area first get "y" in terms of "x" using proportions \[\frac{y}{3-x} = \frac{8}{3}\] \[y = 8-\frac{8}{3}x\] define Area function \[A = x*y = x(8-\frac{8}{3}x)\] take derivative and set equal to 0 \[A' = 8-\frac{16}{3}x = 0\] \[x = \frac{3}{2}\] thus \[y = 4\]

OpenStudy (mony01):

so is 4 the area?

OpenStudy (dumbcow):

wow you didnt follow anything i did huh

OpenStudy (mony01):

to get the area dont you multiply x and y

OpenStudy (dumbcow):

yes correct :)

OpenStudy (mony01):

so the area is 6

OpenStudy (dumbcow):

yep

OpenStudy (mony01):

thank you so much oh and sorry for the first reply...Dont know what I was thinking..

OpenStudy (dumbcow):

yw and its ok, sometimes i get too many stupid replies from other users so it lessens my patience

OpenStudy (dumbcow):

too many times they just look at the last line and assume that must be the answer :{